OpenStudy (iwanttogotostanford):

Find the solution of 3 times the square root of the quantity of x plus 6 equals negative 12, and determine if it is an extraneous solution.

OpenStudy (iwanttogotostanford):

@ospreytriple @BloomLocke367 @freckles

OpenStudy (bloomlocke367):

First off, can you write that in an equation?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (bloomlocke367):

Do that first. Tell us what you get.

OpenStudy (anonymous):

You have\[3\sqrt{x+6} = -12\]What do you think we should do first?

OpenStudy (iwanttogotostanford):

\[3\sqrt{x+6}=-12\]

OpenStudy (iwanttogotostanford):

not sure

OpenStudy (bloomlocke367):

There are a few ways to do this, but the easiest would be to isolate the radical. Do you know how you would do that?

OpenStudy (anonymous):

Well we have to get the 'x' stuff all by itself, so how would you get rid of that 3?

OpenStudy (iwanttogotostanford):

square it?

OpenStudy (bloomlocke367):

you have to get rid of the 3 first. Right now, the 3 is being multiplied with \(\sqrt{x+6}\). How do you "get rid of" something that is being multiplied?

OpenStudy (iwanttogotostanford):

divide

OpenStudy (bloomlocke367):

Yes! now divide both sides by 3 and tell me what you're left with

OpenStudy (iwanttogotostanford):

i don't have a calulator with me right now

OpenStudy (bloomlocke367):

You don't know \(-12\div 3\)?

OpenStudy (iwanttogotostanford):

yes its -4

OpenStudy (bloomlocke367):

There you go. so right now, you have \(\sqrt{x+6}=-4\) correct?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (bloomlocke367):

Now, how would you get rid of that radical?

OpenStudy (iwanttogotostanford):

hmm im not sure

OpenStudy (bloomlocke367):

Well, what's the opposite of taking the square root of something?

OpenStudy (iwanttogotostanford):

squaring it?

OpenStudy (iwanttogotostanford):

im clueless on this sorry

OpenStudy (bloomlocke367):

yes! so you have to square both sides :D

OpenStudy (bloomlocke367):

Don't apologize. You're doing great :D

OpenStudy (iwanttogotostanford):

okay, thanks!

OpenStudy (iwanttogotostanford):

so it would be x^2+36=16?

OpenStudy (bloomlocke367):

No, because when you squared it, it just cancelled out the radical. For instance, when you take \(\sqrt1^2\), you don't get \(1^2\) do you? No, you just get 1. Try again. The right side is correct.

OpenStudy (iwanttogotostanford):

so, x+6=16?

OpenStudy (bloomlocke367):

Yep!

OpenStudy (iwanttogotostanford):

ok, thanks! so then it would be x=10 right?

OpenStudy (iwanttogotostanford):

and would it be extraneous or not extraneous though?

OpenStudy (bloomlocke367):

mhm, I'm pretty certain. Let me ask @Nnesha if I did this correctly. Now we just have to check to make sure it's not extraneous. Do you know what an extraneous solution is?

Nnesha (nnesha):

did a great job!

OpenStudy (bloomlocke367):

Okay, thanks. lol @iwanttogotostanford do you know what an extraneous solution is?

OpenStudy (iwanttogotostanford):

@BloomLocke367 yes

OpenStudy (bloomlocke367):

Then graph it and find out if there's a point where x=4

OpenStudy (iwanttogotostanford):

do you know if it is or not? I'm sorry I'm just in a very hard place right now to be graphing things (I'm on a plane)

Nnesha (nnesha):

you can substitute x for 10 into the original equation to find out if it's extraneous or not. :=)

OpenStudy (iwanttogotostanford):

thanks! @Nnesha

OpenStudy (bloomlocke367):

Sorry, I was afk. Thanks @Nnesha

Nnesha (nnesha):

np

Nnesha (nnesha):

if you get equal sides then it's not extraneous if both sides are not equal then 10 would be an extraneous

OpenStudy (iwanttogotostanford):

ok thanks do u know if it is or not? @Nnesha

OpenStudy (bloomlocke367):

oops, I don't know why I said 4 earlier, I meant to say 10. lol

Nnesha (nnesha):

no i don't. you have to figure it out!

OpenStudy (iwanttogotostanford):

@Nnesha well i don't have paper with me so its really hard

Nnesha (nnesha):

welll, you can use `paint`

Nnesha (nnesha):

that would be fun! trust me

OpenStudy (iwanttogotostanford):

ok, thank you @Nnesha

OpenStudy (iwanttogotostanford):

i got extraneous:-)

Nnesha (nnesha):

both sides are equal ?

OpenStudy (iwanttogotostanford):

no they aren't equal

OpenStudy (bloomlocke367):

There you go!

OpenStudy (iwanttogotostanford):

thanks!

Nnesha (nnesha):

o^_^o

OpenStudy (iwanttogotostanford):

they aren't equal= not extraneous?

OpenStudy (bloomlocke367):

You were correct the first time, it is extraneous because they are not equal.

OpenStudy (bloomlocke367):

wait....

Nnesha (nnesha):

both sides are not equal = extraneous both sides are equal = non-extraneous

OpenStudy (bloomlocke367):

when you take the square root of something, it could be positive or negative

OpenStudy (bloomlocke367):

Right @Nnesha?

Nnesha (nnesha):

`something` if you mean square root of negative number then you will get `imaginary` solution if you take square root of positive then you would always get positive answer

OpenStudy (iwanttogotostanford):

i think its non-extraneous

OpenStudy (bloomlocke367):

how so? because -4*-4=16 AND 4*4=16

OpenStudy (iwanttogotostanford):

oh whoops i did -4 times 4

Nnesha (nnesha):

\[\huge\rm 3\sqrt{10+16}=-12\] whet did you get at left side ?

OpenStudy (iwanttogotostanford):

i get it now

OpenStudy (bloomlocke367):

I gotta go, sorry!

Nnesha (nnesha):

i said `number` if -4 times -4 = positive number so square root would be positive

Nnesha (nnesha):

so it's extraneous or non-extraneous ?

OpenStudy (iwanttogotostanford):

ok it is extraeous then

Nnesha (nnesha):

yes right \[12\cancel{=}-12\]

OpenStudy (iwanttogotostanford):

thank you @Nnesha

Nnesha (nnesha):

np